A fan blade is initially rotating an angular speed of 4.6 rpm. It slows down and eventually comes to rest in a time of 32 seconds after turning through a total of 8.8 revolutions. Find



a) the angular speed in rad/s



b) the average angular velocity and



c) the average angular acceleration of the fan blade.

1
Expert's answer
2022-06-03T07:54:13-0400

"\omega=2\pi\nu=28.9~\frac{rad}s,"

"<\omega>=\frac{\omega+0}2=0.5\omega=14.5~\frac{rad}s,"

"\varepsilon=\frac{\omega^2}{2\varphi}=7.6~\frac{rad}{s^2}."